# A Novel Pathology Foundation Model by Mayo Clinic, Charité, and Aignostics

### v1

@@ Acknowledgments
We would like to thank the teams at Aignostics, Charité - Universitätsmedizin Berlin - Pathology Department, Mayo Clinical Platform, Mayo Clinic Generative Artificial Intelligence Program, Mayo Clinic Department of Laboratory Medicine and Pathology, and Mayo Clinic Digital Pathology for supporting this work. Special thanks goes to Alexander Baxendale, Amelie Froessl, Amin Abbasloo, Barbara Feulner, Carl Andersson, Dharma Indurthy, Erinç Argımak, Fabian Spieß, Gerrit Erdmann, Jay Tolley, Jeannine Korp, Jeff Anderson, Jennifer Flores, Joshua Pankratz, Mark Ibrahim, Olja Smilić, Rob Blundo, Tom Lehmann, Sara Then, Steele Clifton-Berry, and Valentin François for their organizational, technical, and other support.
@@ A.2 Additional Results
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Ours Molecular-related HEST-COAD 25.6 ±2.3$\pm$ 2.3± 2.3 26.2 ±3.1$\pm$ 3.1± 3.1 30.7 ±0.0$\pm$ 0.0± 0.0 31.0 ±2.0$\pm$ 2.0± 2.0 25.9 ±1.6$\pm$ 1.6± 1.6 30.9 ±0.0$\pm$ 0.0± 0.0 29.4 ±1.5$\pm$ 1.5± 1.5 HEST-HCC 7.8 ±1.2$\pm$ 1.2± 1.2 8.3 ±0.5$\pm$ 0.5± 0.5 7.1 ±1.3$\pm$ 1.3± 1.3 9.4 ±1.7$\pm$ 1.7± 1.7 9.6 ±1.0$\pm$ 1.0± 1.0 8.4 ±1.2$\pm$ 1.2± 1.2 10.7 ±1.9$\pm$ 1.9± 1.9 HEST-IDC 56.6 ±7.8$\pm$ 7.8± 7.8 58.5 ±7.7$\pm$ 7.7± 7.7 56.8 ±7.6$\pm$ 7.6± 7.6 57.4 ±8.5$\pm$ 8.5± 8.5 59.3 ±8.5$\pm$ 8.5± 8.5 61.0 ±8.1$\pm$ 8.1± 8.1 60.4 ±8.3$\pm$ 8.3± 8.3 HEST-LUAD 54.8 ±2.3$\pm$ 2.3± 2.3 55.2 ±2.2$\pm$ 2.2± 2.2 55.8 ±2.9$\pm$ 2.9± 2.9 57.7 ±1.8$\pm$ 1.8± 1.8 56.9 ±1.7$\pm$ 1.7± 1.7 57.3 ±2.7$\pm$ 2.7± 2.7 58.0 ±1.5$\pm$ 1.5± 1.5 HEST-LYMPH_IDC 24.8 ±4.9$\pm$ 4.9± 4.9 25.8 ±4.1$\pm$ 4.1± 4.1 25.1 ±4.2$\pm$ 4.2± 4.2 25.6 ±3.3$\pm$ 3.3± 3.3 25.9 ±3.3$\pm$ 3.3± 3.3 26.8 ±4.0$\pm$ 4.0± 4.0 26.4 ±4.4$\pm$ 4.4± 4.4 HEST-PAAD 47.9 ±7.0$\pm$ 7.0± 7.0 48.8 ±5.8$\pm$ 5.8± 5.8 49.5 ±5.8$\pm$ 5.8± 5.8 51.1 ±7.8$\pm$ 7.8± 7.8 47.3 ±6.9$\pm$ 6.9± 6.9 50.9 ±4.3$\pm$ 4.3± 4.3 51.8 ±7.4$\pm$ 7.4± 7.4 HEST-PRAD 37.7 ±0.1$\pm$ 0.1± 0.1 32.2 ±8.1$\pm$ 8.1± 8.1 38.4 ±3.1$\pm$ 3.1± 3.1 37.7 ±2.6$\pm$ 2.6± 2.6 35.1 ±3.4$\pm$ 3.4± 3.4 38.5 ±0.0$\pm$ 0.0± 0.0 38.4 ±0.7$\pm$ 0.7± 0.7 HEST-READ 18.5 ±5.9$\pm$ 5.9± 5.9 18.4 ±4.9$\pm$ 4.9± 4.9 19.6 ±6.2$\pm$ 6.2± 6.2 19.9 ±6.8$\pm$ 6.8± 6.8 21.1 ±5.0$\pm$ 5.0± 5.0 24.1 ±2.7$\pm$ 2.7± 2.7 22.8 ±3.1$\pm$ 3.1± 3.1 HEST-SKCM 58.4 ±6.2$\pm$ 6.2± 6.2 63.5 ±3.6$\pm$ 3.6± 3.6 58.8 ±5.8$\pm$ 5.8± 5.8 61.8 ±4.6$\pm$ 4.6± 4.6 63.7 ±3.1$\pm$ 3.1± 3.1 66.1 ±5.8$\pm$ 5.8± 5.8 62.5 ±2.4$\pm$ 2.4± 2.4 HEST-ccRCC 27.3 ±4.0$\pm$ 4.0± 4.0 25.3 ±3.8$\pm$ 3.8± 3.8 24.9 ±4.1$\pm$ 4.1± 4.1 25.3 ±5.2$\pm$ 5.2± 5.2 27.4 ±4.5$\pm$ 4.5± 4.5 29.0 ±3.8$\pm$ 3.8± 3.8 29.4 ±5.5$\pm$ 5.5± 5.5 MSI CRC (patch) 68.8 ±0.1$\pm$ 0.1± 0.1 69.5 ±0.0$\pm$ 0.0± 0.0 70.4 ±0.1$\pm$ 0.1± 0.1 69.9 ±0.1$\pm$ 0.1± 0.1 74.0 ±0.0$\pm$ 0.0± 0.0 71.2 ±0.1$\pm$ 0.1± 0.1 73.6 ±0.0$\pm$ 0.0± 0.0 MSI STAD (patch) 71.2 ±0.1$\pm$ 0.1± 0.1 70.5 ±0.0$\pm$ 0.0± 0.0 71.0 ±0.1$\pm$ 0.1± 0.1 74.1 ±0.1$\pm$ 0.1± 0.1 74.8 ±0.2$\pm$ 0.2± 0.2 73.6 ±0.0$\pm$ 0.0± 0.0 76.0 ±0.1$\pm$ 0.1± 0.1 Molecular-Average 41.6 41.8 42.3 43.4 43.4 44.8 44.9 Morphology-related Pan-cancer TIL 92.9 ±0.0$\pm$ 0.0± 0.0 92.6 ±0.0$\pm$ 0.0± 0.0 92.3 ±0.0$\pm$ 0.0± 0.0 92.6 ±0.1$\pm$ 0.1± 0.1 93.1 ±0.1$\pm$ 0.1± 0.1 93.0 ±0.1$\pm$ 0.1± 0.1 93.0 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (10x) 64.0 ±0.0$\pm$ 0.0± 0.0 68.6 ±0.0$\pm$ 0.0± 0.0 69.1 ±0.0$\pm$ 0.0± 0.0 70.6 ±0.0$\pm$ 0.0± 0.0 73.0 ±0.0$\pm$ 0.0± 0.0 70.4 ±0.0$\pm$ 0.0± 0.0 71.8 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (20x) 69.8 ±0.0$\pm$ 0.0± 0.0 67.8 ±0.0$\pm$ 0.0± 0.0 68.0 ±0.0$\pm$ 0.0± 0.0 78.1 ±0.0$\pm$ 0.0± 0.0 71.5 ±0.0$\pm$ 0.0± 0.0 72.4 ±0.0$\pm$ 0.0± 0.0 67.8 ±0.0$\pm$ 0.0± 0.0 BACH 73.8 ±0.3$\pm$ 0.3± 0.3 80.1 ±1.0$\pm$ 1.0± 1.0 80.2 ±0.3$\pm$ 0.3± 0.3 76.9 ±0.3$\pm$ 0.3± 0.3 88.7 ±0.3$\pm$ 0.3± 0.3 75.8 ±1.1$\pm$ 1.1± 1.1 93.1 ±0.2$\pm$ 0.2± 0.2 CRC-100k 95.5 ±0.0$\pm$ 0.0± 0.0 95.4 ±0.2$\pm$ 0.2± 0.2 95.9 ±0.0$\pm$ 0.0± 0.0 96.0 ±0.1$\pm$ 0.1± 0.1 96.7 ±0.1$\pm$ 0.1± 0.1 96.2 ±0.1$\pm$ 0.1± 0.1 97.1 ±0.1$\pm$ 0.1± 0.1 MHIST 78.4 ±0.4$\pm$ 0.4± 0.4 84.4 ±0.1$\pm$ 0.1± 0.1 83.1 ±0.1$\pm$ 0.1± 0.1 80.5 ±0.0$\pm$ 0.0± 0.0 85.9 ±0.1$\pm$ 0.1± 0.1 85.0 ±0.2$\pm$ 0.2± 0.2 86.4 ±0.1$\pm$ 0.1± 0.1 PCAM 90.0 ±0.1$\pm$ 0.1± 0.1 93.6 ±0.1$\pm$ 0.1± 0.1 94.5 ±0.0$\pm$ 0.0± 0.0 94.6 ±0.1$\pm$ 0.1± 0.1 93.9 ±0.1$\pm$ 0.1± 0.1 94.3 ±0.1$\pm$ 0.1± 0.1 94.9 ±0.0$\pm$ 0.0± 0.0 CAMELYON16 79.8 ±1.4$\pm$ 1.4± 1.4 85.0 ±0.5$\pm$ 0.5± 0.5 82.1 ±0.7$\pm$ 0.7± 0.7 77.1 ±1.7$\pm$ 1.7± 1.7 86.5 ±0.2$\pm$ 0.2± 0.2 84.0 ±0.8$\pm$ 0.8± 0.8 86.8 ±0.4$\pm$ 0.4± 0.4 PANDA 65.3 ±0.4$\pm$ 0.4± 0.4 69.6 ±0.6$\pm$ 0.6± 0.6 69.6 ±0.7$\pm$ 0.7± 0.7 69.6 ±0.4$\pm$ 0.4± 0.4 66.4 ±1.1$\pm$ 1.1± 1.1 68.0 ±0.6$\pm$ 0.6± 0.6 70.5 ±0.5$\pm$ 0.5± 0.5 Morphology-Average 78.8 81.9 81.6 81.8 84.0 82.1 84.6 Overall Average 57.6 59.0 59.2 59.9 60.8 60.8 61.9
@@
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Ours Molecular-related HEST-COAD 25.0 ±1.7$\pm$ 1.7± 1.7 26.2 ±3.1$\pm$ 3.1± 3.1 29.9 ±2.1$\pm$ 2.1± 2.1 23.7 ±6.0$\pm$ 6.0± 6.0 25.9 ±1.6$\pm$ 1.6± 1.6 30.9 ±0.0$\pm$ 0.0± 0.0 25.9 ±3.1$\pm$ 3.1± 3.1 HEST-HCC 6.7 ±1.3$\pm$ 1.3± 1.3 7.8 ±0.2$\pm$ 0.2± 0.2 7.1 ±1.3$\pm$ 1.3± 1.3 6.5 ±0.1$\pm$ 0.1± 0.1 8.2 ±1.0$\pm$ 1.0± 1.0 7.9 ±0.6$\pm$ 0.6± 0.6 9.4 ±0.8$\pm$ 0.8± 0.8 HEST-IDC 54.1 ±7.7$\pm$ 7.7± 7.7 57.4 ±7.9$\pm$ 7.9± 7.9 55.1 ±7.3$\pm$ 7.3± 7.3 54.5 ±8.9$\pm$ 8.9± 8.9 59.2 ±8.0$\pm$ 8.0± 8.0 59.8 ±8.5$\pm$ 8.5± 8.5 59.6 ±8.1$\pm$ 8.1± 8.1 HEST-LUAD 54.2 ±1.1$\pm$ 1.1± 1.1 54.6 ±2.2$\pm$ 2.2± 2.2 54.1 ±3.6$\pm$ 3.6± 3.6 55.5 ±1.2$\pm$ 1.2± 1.2 55.3 ±1.7$\pm$ 1.7± 1.7 55.9 ±3.3$\pm$ 3.3± 3.3 57.0 ±1.7$\pm$ 1.7± 1.7 HEST-LYMPH_IDC 24.4 ±4.6$\pm$ 4.6± 4.6 25.6 ±4.4$\pm$ 4.4± 4.4 25.0 ±5.0$\pm$ 5.0± 5.0 24.3 ±3.5$\pm$ 3.5± 3.5 25.5 ±2.6$\pm$ 2.6± 2.6 25.9 ±4.0$\pm$ 4.0± 4.0 25.7 ±4.7$\pm$ 4.7± 4.7 HEST-PAAD 44.5 ±6.6$\pm$ 6.6± 6.6 48.1 ±7.0$\pm$ 7.0± 7.0 47.5 ±4.8$\pm$ 4.8± 4.8 45.7 ±5.8$\pm$ 5.8± 5.8 47.2 ±6.5$\pm$ 6.5± 6.5 49.1 ±4.0$\pm$ 4.0± 4.0 50.7 ±7.2$\pm$ 7.2± 7.2 HEST-PRAD 35.4 ±1.5$\pm$ 1.5± 1.5 29.4 ±8.5$\pm$ 8.5± 8.5 37.0 ±2.1$\pm$ 2.1± 2.1 37.0 ±2.8$\pm$ 2.8± 2.8 34.8 ±3.1$\pm$ 3.1± 3.1 38.5 ±0.0$\pm$ 0.0± 0.0 35.3 ±3.2$\pm$ 3.2± 3.2 HEST-READ 17.5 ±5.9$\pm$ 5.9± 5.9 18.4 ±4.9$\pm$ 4.9± 4.9 19.6 ±6.2$\pm$ 6.2± 6.2 17.6 ±8.1$\pm$ 8.1± 8.1 20.9 ±5.0$\pm$ 5.0± 5.0 22.2 ±4.8$\pm$ 4.8± 4.8 21.3 ±2.9$\pm$ 2.9± 2.9 HEST-SKCM 55.5 ±3.6$\pm$ 3.6± 3.6 63.5 ±3.6$\pm$ 3.6± 3.6 56.1 ±6.2$\pm$ 6.2± 6.2 58.0 ±4.1$\pm$ 4.1± 4.1 61.9 ±2.8$\pm$ 2.8± 2.8 64.5 ±6.2$\pm$ 6.2± 6.2 56.2 ±0.5$\pm$ 0.5± 0.5 HEST-ccRCC 26.6 ±3.8$\pm$ 3.8± 3.8 24.0 ±4.0$\pm$ 4.0± 4.0 24.3 ±3.3$\pm$ 3.3± 3.3 24.9 ±5.4$\pm$ 5.4± 5.4 27.4 ±4.5$\pm$ 4.5± 4.5 26.8 ±3.2$\pm$ 3.2± 3.2 27.8 ±3.6$\pm$ 3.6± 3.6 MSI CRC (patch) 67.5 ±0.0$\pm$ 0.0± 0.0 69.1 ±0.0$\pm$ 0.0± 0.0 69.0 ±0.1$\pm$ 0.1± 0.1 68.0 ±0.0$\pm$ 0.0± 0.0 71.6 ±0.1$\pm$ 0.1± 0.1 69.7 ±0.0$\pm$ 0.0± 0.0 71.6 ±0.0$\pm$ 0.0± 0.0 MSI STAD (patch) 68.6 ±0.0$\pm$ 0.0± 0.0 68.6 ±0.0$\pm$ 0.0± 0.0 67.4 ±0.1$\pm$ 0.1± 0.1 72.4 ±0.0$\pm$ 0.0± 0.0 72.8 ±0.0$\pm$ 0.0± 0.0 72.7 ±0.0$\pm$ 0.0± 0.0 73.5 ±0.0$\pm$ 0.0± 0.0 Molecular-Average 40.0 41.1 41.0 40.7 42.6 43.6 42.8 Morphology-related Pan-cancer TIL 92.6 ±0.0$\pm$ 0.0± 0.0 92.4 ±0.0$\pm$ 0.0± 0.0 91.8 ±0.0$\pm$ 0.0± 0.0 91.9 ±0.0$\pm$ 0.0± 0.0 92.7 ±0.0$\pm$ 0.0± 0.0 92.6 ±0.0$\pm$ 0.0± 0.0 92.8 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (10x) 63.9 ±0.0$\pm$ 0.0± 0.0 68.3 ±0.0$\pm$ 0.0± 0.0 68.7 ±0.0$\pm$ 0.0± 0.0 70.2 ±0.0$\pm$ 0.0± 0.0 72.9 ±0.0$\pm$ 0.0± 0.0 69.9 ±0.0$\pm$ 0.0± 0.0 71.8 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (20x) 69.8 ±0.0$\pm$ 0.0± 0.0 67.4 ±0.0$\pm$ 0.0± 0.0 67.4 ±0.0$\pm$ 0.0± 0.0 77.7 ±0.0$\pm$ 0.0± 0.0 71.5 ±0.0$\pm$ 0.0± 0.0 72.1 ±0.0$\pm$ 0.0± 0.0 67.2 ±0.0$\pm$ 0.0± 0.0 BACH 72.7 ±0.3$\pm$ 0.3± 0.3 79.7 ±0.4$\pm$ 0.4± 0.4 76.1 ±0.4$\pm$ 0.4± 0.4 74.9 ±0.5$\pm$ 0.5± 0.5 88.0 ±0.4$\pm$ 0.4± 0.4 75.8 ±1.1$\pm$ 1.1± 1.1 93.1 ±0.2$\pm$ 0.2± 0.2 CRC-100k 93.9 ±0.0$\pm$ 0.0± 0.0 94.8 ±0.1$\pm$ 0.1± 0.1 95.2 ±0.1$\pm$ 0.1± 0.1 94.8 ±0.1$\pm$ 0.1± 0.1 96.6 ±0.1$\pm$ 0.1± 0.1 95.8 ±0.1$\pm$ 0.1± 0.1 97.0 ±0.0$\pm$ 0.0± 0.0 MHIST 77.5 ±0.1$\pm$ 0.1± 0.1 84.4 ±0.1$\pm$ 0.1± 0.1 82.9 ±0.1$\pm$ 0.1± 0.1 79.8 ±0.1$\pm$ 0.1± 0.1 85.8 ±0.2$\pm$ 0.2± 0.2 83.9 ±0.1$\pm$ 0.1± 0.1 85.2 ±0.1$\pm$ 0.1± 0.1 PCAM 89.3 ±0.0$\pm$ 0.0± 0.0 93.6 ±0.1$\pm$ 0.1± 0.1 94.5 ±0.0$\pm$ 0.0± 0.0 94.2 ±0.1$\pm$ 0.1± 0.1 93.6 ±0.1$\pm$ 0.1± 0.1 94.2 ±0.1$\pm$ 0.1± 0.1 94.9 ±0.0$\pm$ 0.0± 0.0 CAMELYON16 79.8 ±1.4$\pm$ 1.4± 1.4 84.9 ±0.8$\pm$ 0.8± 0.8 82.1 ±0.7$\pm$ 0.7± 0.7 77.1 ±1.7$\pm$ 1.7± 1.7 86.5 ±0.2$\pm$ 0.2± 0.2 83.2 ±1.4$\pm$ 1.4± 1.4 86.8 ±0.4$\pm$ 0.4± 0.4 PANDA 64.3 ±0.4$\pm$ 0.4± 0.4 69.6 ±0.6$\pm$ 0.6± 0.6 66.1 ±0.3$\pm$ 0.3± 0.3 68.2 ±0.7$\pm$ 0.7± 0.7 65.1 ±0.9$\pm$ 0.9± 0.9 67.2 ±0.4$\pm$ 0.4± 0.4 70.0 ±0.5$\pm$ 0.5± 0.5 Morphology-Average 78.2 81.7 80.5 81.0 83.6 81.6 84.3 Overall Average 56.4 58.5 57.9 57.9 60.2 59.9 60.6

### v2

@@ Acknowledgments
We would like to thank the teams at Aignostics, Charité - Universitätsmedizin Berlin - Pathology Department, Mayo Clinical Platform, Mayo Clinic Generative Artificial Intelligence Program, Mayo Clinic Department of Laboratory Medicine and Pathology, and Mayo Clinic Digital Pathology for supporting this work. Special thanks goes to Alexander Baxendale, Amelie Froessl, Amin Abbasloo, Barbara Feulner, Carl Andersson, Dharma Indurthy, Erinç Argımak, Fabian Spieß, Gerrit Erdmann, Jay Tolley, Jeannine Korp, Jeff Anderson, Jennifer Flores, Joshua Pankratz, Julika Ribbat-Idel, Mark Ibrahim, Olja Smilić, Rob Blundo, Tom Lehmann, Sara Then, Steele Clifton-Berry, and Valentin François for their organizational, technical, and other support.
@@ A.2 Additional Results
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Atlas Molecular-related HEST-COAD 25.6 ±$\pm 2.3$ 26.2 ±$\pm 3.1$ 30.7 ±$\pm 0.0$ 31.0 ±$\pm 2.0$ 25.9 ±$\pm 1.6$ 30.9 ±$\pm 0.0$ 29.4 ±$\pm 1.5$ HEST-HCC 7.8 ±$\pm 1.2$ 8.3 ±$\pm 0.5$ 7.1 ±$\pm 1.3$ 9.4 ±$\pm 1.7$ 9.6 ±$\pm 1.0$ 8.4 ±$\pm 1.2$ 10.7 ±$\pm 1.9$ HEST-IDC 56.6 ±$\pm 7.8$ 58.5 ±$\pm 7.7$ 56.8 ±$\pm 7.6$ 57.4 ±$\pm 8.5$ 59.3 ±$\pm 8.5$ 61.0 ±$\pm 8.1$ 60.4 ±$\pm 8.3$ HEST-LUAD 54.8 ±$\pm 2.3$ 55.2 ±$\pm 2.2$ 55.8 ±$\pm 2.9$ 57.7 ±$\pm 1.8$ 56.9 ±$\pm 1.7$ 57.3 ±$\pm 2.7$ 58.0 ±$\pm 1.5$ HEST-LYMPH_IDC 24.8 ±$\pm 4.9$ 25.8 ±$\pm 4.1$ 25.1 ±$\pm 4.2$ 25.6 ±$\pm 3.3$ 25.9 ±$\pm 3.3$ 26.8 ±$\pm 4.0$ 26.4 ±$\pm 4.4$ HEST-PAAD 47.9 ±$\pm 7.0$ 48.8 ±$\pm 5.8$ 49.5 ±$\pm 5.8$ 51.1 ±$\pm 7.8$ 47.3 ±$\pm 6.9$ 50.9 ±$\pm 4.3$ 51.8 ±$\pm 7.4$ HEST-PRAD 37.7 ±$\pm 0.1$ 32.2 ±$\pm 8.1$ 38.4 ±$\pm 3.1$ 37.7 ±$\pm 2.6$ 35.1 ±$\pm 3.4$ 38.5 ±$\pm 0.0$ 38.4 ±$\pm 0.7$ HEST-READ 18.5 ±$\pm 5.9$ 18.4 ±$\pm 4.9$ 19.6 ±$\pm 6.2$ 19.9 ±$\pm 6.8$ 21.1 ±$\pm 5.0$ 24.1 ±$\pm 2.7$ 22.8 ±$\pm 3.1$ HEST-SKCM 58.4 ±$\pm 6.2$ 63.5 ±$\pm 3.6$ 58.8 ±$\pm 5.8$ 61.8 ±$\pm 4.6$ 63.7 ±$\pm 3.1$ 66.1 ±$\pm 5.8$ 62.5 ±$\pm 2.4$ HEST-ccRCC 27.3 ±$\pm 4.0$ 25.3 ±$\pm 3.8$ 24.9 ±$\pm 4.1$ 25.3 ±$\pm 5.2$ 27.4 ±$\pm 4.5$ 29.0 ±$\pm 3.8$ 29.4 ±$\pm 5.5$ MSI CRC (patch) 68.8 ±$\pm 0.1$ 69.5 ±$\pm 0.0$ 70.4 ±$\pm 0.1$ 69.9 ±$\pm 0.1$ 74.0 ±$\pm 0.0$ 71.2 ±$\pm 0.1$ 73.6 ±$\pm 0.0$ MSI STAD (patch) 71.2 ±$\pm 0.1$ 70.5 ±$\pm 0.0$ 71.0 ±$\pm 0.1$ 74.1 ±$\pm 0.1$ 74.8 ±$\pm 0.2$ 73.6 ±$\pm 0.0$ 76.0 ±$\pm 0.1$ Molecular-Average 41.6 41.8 42.3 43.4 43.4 44.8 44.9 Morphology-related Pan-cancer TIL 92.9 ±$\pm 0.0$ 92.6 ±$\pm 0.0$ 92.3 ±$\pm 0.0$ 92.6 ±$\pm 0.1$ 93.1 ±$\pm 0.1$ 93.0 ±$\pm 0.1$ 93.0 ±$\pm 0.0$ TCGA Uniform (10x) 64.0 ±$\pm 0.0$ 68.6 ±$\pm 0.0$ 69.1 ±$\pm 0.0$ 70.6 ±$\pm 0.0$ 73.0 ±$\pm 0.0$ 70.4 ±$\pm 0.0$ 71.8 ±$\pm 0.0$ TCGA Uniform (20x) 69.8 ±$\pm 0.0$ 67.8 ±$\pm 0.0$ 68.0 ±$\pm 0.0$ 78.1 ±$\pm 0.0$ 71.5 ±$\pm 0.0$ 72.4 ±$\pm 0.0$ 67.8 ±$\pm 0.0$ BACH 73.8 ±$\pm 0.3$ 80.1 ±$\pm 1.0$ 80.2 ±$\pm 0.3$ 76.9 ±$\pm 0.3$ 88.7 ±$\pm 0.3$ 75.8 ±$\pm 1.1$ 93.1 ±$\pm 0.2$ CRC-100k 95.5 ±$\pm 0.0$ 95.4 ±$\pm 0.2$ 95.9 ±$\pm 0.0$ 96.0 ±$\pm 0.1$ 96.7 ±$\pm 0.1$ 96.2 ±$\pm 0.1$ 97.1 ±$\pm 0.1$ MHIST 78.4 ±$\pm 0.4$ 84.4 ±$\pm 0.1$ 83.1 ±$\pm 0.1$ 80.5 ±$\pm 0.0$ 85.9 ±$\pm 0.1$ 85.0 ±$\pm 0.2$ 86.4 ±$\pm 0.1$ PCAM 90.0 ±$\pm 0.1$ 93.6 ±$\pm 0.1$ 94.5 ±$\pm 0.0$ 94.6 ±$\pm 0.1$ 93.9 ±$\pm 0.1$ 94.3 ±$\pm 0.1$ 94.9 ±$\pm 0.0$ CAMELYON16 79.8 ±$\pm 1.4$ 85.0 ±$\pm 0.5$ 82.1 ±$\pm 0.7$ 77.1 ±$\pm 1.7$ 86.5 ±$\pm 0.2$ 84.0 ±$\pm 0.8$ 86.8 ±$\pm 0.4$ PANDA 65.3 ±$\pm 0.4$ 69.6 ±$\pm 0.6$ 69.6 ±$\pm 0.7$ 69.6 ±$\pm 0.4$ 66.4 ±$\pm 1.1$ 68.0 ±$\pm 0.6$ 70.5 ±$\pm 0.5$ Morphology-Average 78.8 81.9 81.6 81.8 84.0 82.1 84.6 Overall Average 57.6 59.0 59.2 59.9 60.8 60.8 61.9
@@
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Atlas Molecular-related HEST-COAD 25.0 ±$\pm 1.7$ 26.2 ±$\pm 3.1$ 29.9 ±$\pm 2.1$ 23.7 ±$\pm 6.0$ 25.9 ±$\pm 1.6$ 30.9 ±$\pm 0.0$ 25.9 ±$\pm 3.1$ HEST-HCC 6.7 ±$\pm 1.3$ 7.8 ±$\pm 0.2$ 7.1 ±$\pm 1.3$ 6.5 ±$\pm 0.1$ 8.2 ±$\pm 1.0$ 7.9 ±$\pm 0.6$ 9.4 ±$\pm 0.8$ HEST-IDC 54.1 ±$\pm 7.7$ 57.4 ±$\pm 7.9$ 55.1 ±$\pm 7.3$ 54.5 ±$\pm 8.9$ 59.2 ±$\pm 8.0$ 59.8 ±$\pm 8.5$ 59.6 ±$\pm 8.1$ HEST-LUAD 54.2 ±$\pm 1.1$ 54.6 ±$\pm 2.2$ 54.1 ±$\pm 3.6$ 55.5 ±$\pm 1.2$ 55.3 ±$\pm 1.7$ 55.9 ±$\pm 3.3$ 57.0 ±$\pm 1.7$ HEST-LYMPH_IDC 24.4 ±$\pm 4.6$ 25.6 ±$\pm 4.4$ 25.0 ±$\pm 5.0$ 24.3 ±$\pm 3.5$ 25.5 ±$\pm 2.6$ 25.9 ±$\pm 4.0$ 25.7 ±$\pm 4.7$ HEST-PAAD 44.5 ±$\pm 6.6$ 48.1 ±$\pm 7.0$ 47.5 ±$\pm 4.8$ 45.7 ±$\pm 5.8$ 47.2 ±$\pm 6.5$ 49.1 ±$\pm 4.0$ 50.7 ±$\pm 7.2$ HEST-PRAD 35.4 ±$\pm 1.5$ 29.4 ±$\pm 8.5$ 37.0 ±$\pm 2.1$ 37.0 ±$\pm 2.8$ 34.8 ±$\pm 3.1$ 38.5 ±$\pm 0.0$ 35.3 ±$\pm 3.2$ HEST-READ 17.5 ±$\pm 5.9$ 18.4 ±$\pm 4.9$ 19.6 ±$\pm 6.2$ 17.6 ±$\pm 8.1$ 20.9 ±$\pm 5.0$ 22.2 ±$\pm 4.8$ 21.3 ±$\pm 2.9$ HEST-SKCM 55.5 ±$\pm 3.6$ 63.5 ±$\pm 3.6$ 56.1 ±$\pm 6.2$ 58.0 ±$\pm 4.1$ 61.9 ±$\pm 2.8$ 64.5 ±$\pm 6.2$ 56.2 ±$\pm 0.5$ HEST-ccRCC 26.6 ±$\pm 3.8$ 24.0 ±$\pm 4.0$ 24.3 ±$\pm 3.3$ 24.9 ±$\pm 5.4$ 27.4 ±$\pm 4.5$ 26.8 ±$\pm 3.2$ 27.8 ±$\pm 3.6$ MSI CRC (patch) 67.5 ±$\pm 0.0$ 69.1 ±$\pm 0.0$ 69.0 ±$\pm 0.1$ 68.0 ±$\pm 0.0$ 71.6 ±$\pm 0.1$ 69.7 ±$\pm 0.0$ 71.6 ±$\pm 0.0$ MSI STAD (patch) 68.6 ±$\pm 0.0$ 68.6 ±$\pm 0.0$ 67.4 ±$\pm 0.1$ 72.4 ±$\pm 0.0$ 72.8 ±$\pm 0.0$ 72.7 ±$\pm 0.0$ 73.5 ±$\pm 0.0$ Molecular-Average 40.0 41.1 41.0 40.7 42.6 43.6 42.8 Morphology-related Pan-cancer TIL 92.6 ±$\pm 0.0$ 92.4 ±$\pm 0.0$ 91.8 ±$\pm 0.0$ 91.9 ±$\pm 0.0$ 92.7 ±$\pm 0.0$ 92.6 ±$\pm 0.0$ 92.8 ±$\pm 0.0$ TCGA Uniform (10x) 63.9 ±$\pm 0.0$ 68.3 ±$\pm 0.0$ 68.7 ±$\pm 0.0$ 70.2 ±$\pm 0.0$ 72.9 ±$\pm 0.0$ 69.9 ±$\pm 0.0$ 71.8 ±$\pm 0.0$ TCGA Uniform (20x) 69.8 ±$\pm 0.0$ 67.4 ±$\pm 0.0$ 67.4 ±$\pm 0.0$ 77.7 ±$\pm 0.0$ 71.5 ±$\pm 0.0$ 72.1 ±$\pm 0.0$ 67.2 ±$\pm 0.0$ BACH 72.7 ±$\pm 0.3$ 79.7 ±$\pm 0.4$ 76.1 ±$\pm 0.4$ 74.9 ±$\pm 0.5$ 88.0 ±$\pm 0.4$ 75.8 ±$\pm 1.1$ 93.1 ±$\pm 0.2$ CRC-100k 93.9 ±$\pm 0.0$ 94.8 ±$\pm 0.1$ 95.2 ±$\pm 0.1$ 94.8 ±$\pm 0.1$ 96.6 ±$\pm 0.1$ 95.8 ±$\pm 0.1$ 97.0 ±$\pm 0.0$ MHIST 77.5 ±$\pm 0.1$ 84.4 ±$\pm 0.1$ 82.9 ±$\pm 0.1$ 79.8 ±$\pm 0.1$ 85.8 ±$\pm 0.2$ 83.9 ±$\pm 0.1$ 85.2 ±$\pm 0.1$ PCAM 89.3 ±$\pm 0.0$ 93.6 ±$\pm 0.1$ 94.5 ±$\pm 0.0$ 94.2 ±$\pm 0.1$ 93.6 ±$\pm 0.1$ 94.2 ±$\pm 0.1$ 94.9 ±$\pm 0.0$ CAMELYON16 79.8 ±$\pm 1.4$ 84.9 ±$\pm 0.8$ 82.1 ±$\pm 0.7$ 77.1 ±$\pm 1.7$ 86.5 ±$\pm 0.2$ 83.2 ±$\pm 1.4$ 86.8 ±$\pm 0.4$ PANDA 64.3 ±$\pm 0.4$ 69.6 ±$\pm 0.6$ 66.1 ±$\pm 0.3$ 68.2 ±$\pm 0.7$ 65.1 ±$\pm 0.9$ 67.2 ±$\pm 0.4$ 70.0 ±$\pm 0.5$ Morphology-Average 78.2 81.7 80.5 81.0 83.6 81.6 84.3 Overall Average 56.4 58.5 57.9 57.9 60.2 59.9 60.6